# Localizing Factual Inconsistencies in Attributable Text Generation

### v1

@@ Abstract
There has been an increasing interest in detecting hallucinations in model-generated texts, both manually and automatically, at varying levels of granularity. However, most existing methods fail to precisely pinpoint the errors. In this work, we introduce QASemConsistency, a new formalism for localizing factual inconsistencies in attributable text generation, at a fine-grained level. Drawing inspiration from Neo-Davidsonian formal semantics, we propose decomposing the generated text into minimal predicate-argument level propositions, expressed as simple question-answer (QA) pairs, and assess whether each individual QA pair is supported by a trusted reference text. As each QA pair corresponds to a single semantic relation between a predicate and an argument, QASemConsistency effectively localizes the unsupported information. We first demonstrate the effectiveness of the QASemConsistency methodology for human annotation, by collecting crowdsourced annotations of granular consistency errors, while achieving a substantial inter-annotator agreement ($\kappa>0.7)$. Then, we implement several methods for automatically detecting localized factual inconsistencies, with both supervised entailment models and open-source LLMs. [^1] Our codebase, dataset, and models can be found at https://github.com/ariecattan/qasem_consistency
@@ 1 Introduction
Figure 1: An annotation example of a localized factual consistency error according to our QASemConsistency methodology. Here, the model incorrectly identifies the location of the paraglider’s fall as the location where the rescue team was training, indicated by the QA “Where did someone fall? in the Annalong Valley in Countly Antrim”.
@@
This work focuses on attributable text generation, where the generated content can be verified against a trusted supporting source, referred here as the “reference text”. This reference text may be part of the input for generation, as typical in text summarization or open book QA (Gao et al., 2023b; Slobodkin et al., 2024), retrieved post-generation (Bohnet et al., 2022; Gao et al., 2023a; Min et al., 2023), or identified by the models themselves when instructed to provide citations to external sources (Liu et al., 2023a; Yue et al., 2023).
@@
In this work, we introduce QASemConsistency, a novel protocol for detecting localized factual inconsistencies in attributable text generation, applicable to both human and automatic detection. Inspired by Neo-Davidsonian formal semantics, our method decomposes the generated text into elementary assertions, in the form of atomic question-answer (QA) pairs (QASRL (He et al., 2015) and QANom (Klein et al., 2020)), where each pair corresponds to a single predicate-argument relation. Localizing factual inconsistencies then involves identifying the set of QA pairs that are not supported by the reference text. Figure 1 illustrates our QASemConsistency methodology by representing the factual inconsistency via a simple QA “Where did someone fall? in the Annalong Valley in County Antrim”, which is not supported by the reference text. By assessing each individual predicate-argument level statement, QASemConsistency can pinpoint more precisely to the factual mistakes than prior methods (illustrated in Table 1).
Furthermore, since we represent predicate-argument relations with simple natural language expressions (questions and answers), our QASemConsistency methodology is well-suited for manually annotating granular consistency errors. Indeed, we collect the first dataset with localized annotations of factual inconsistency at the predicate-argument level via crowdsourcing and achieve a high inter-annotator agreement (> 0.7 Fleiss Kappa) (§4).
Finally, we implement methods for automatically detecting whether each individual QASem QA is supported by the reference text, following the QASemConsistency methodology (§5). We conduct experiments with a variety of models, including supervised NLI models and prompting several LLMs. While these models were supervised on standard entailment datasets, our results show that they can effectively handle our fine-grained QA annotation, providing more detailed error detection.
@@ 2.2 Predicate-argument Relations: From Neo-Davidsonian Semantics to QA-SRL
{dmath}
∃e_1, e_2   (Buying(e_1) ∧Agent(e_1, Mary) ∧Theme(e_1, Book) ∧Time(e_1, Yesterday) ∧Giving(e_2) ∧Agent(e_2, Mary) ∧Recipient(e_2, John) ∧Theme(e_2, Book) ∧Manner(e_2, Smile))
@@
To bridge this gap, we propose using QA-SRL, a semantic formalism that simplifies over traditional SRL schemes by representing each predicate-argument relation through a simple and “minimal” question-answer pair, such as “Who bought something? Mary” (He et al., 2015; FitzGerald et al., 2018; Roit et al., 2020). For example, the QA-SRL representation of the sentence above is detailed in 2. Here, “Mary” is identified as the agent of the predicate “bought”, where this semantic relation is represented by the question “Who bought something? Mary”. Loosely speaking, each QA typically corresponds to a single Neo-Davidsonian proposition.
@@
By relying on a comprehensible natural-language representation, QA-SRL largely subsumes traditional SRL schemes. Notably, it successfully covers valuable implicit semantic arguments (Roit et al., 2024), which are intuitively captured by human annotators, and subsequently by models trained on such annotated data. The QA-SRL formalism has been expanded to include discourse relations (e.g., QADiscourse; Pyatkin et al., 2020), deverbal nominalizations (e.g., QANom; Klein et al., 2020) and adjectival predicates  (QA-Adj; Pesahov et al., 2023), along with the development of a unified QASem framework, covering multiple predicate types (Klein et al., 2022).
@@ 3 QASemConsistency
Given a generated text $\mathcal{Y}$ and a reference text $\mathcal{X}$ that is expected to support the information in $\mathcal{Y}$ (e.g. the source for generation or a grounding text for it), we define the task of localizing factual inconsistencies as identifying the set of “localized” assertions that are contained in $\mathcal{Y}$ but are not supported by $\mathcal{X}$. This involves the decomposition of $\mathcal{Y}$ into such assertions, each being a unit of information that can be individually assessed for its entailment by $\mathcal{X}$. For effective localization, we suggest two desired properties of this decomposition. First, the assertions should be as minimal in scope as possible, where ideally each assertion should not be further decomposable into smaller verifiable assertions. Second, each assertion should be human interpretable, providing clear insights to common language speakers, thus allowing efficient crowdsourced annotation for localized factual inconsistencies.
To fulfill these properties, we propose decomposing $\mathcal{Y}$ into the set of its predicate-argument level propositions, using the QASem framework (§2.2). By construction, each question-answer pair (QA) in QASem corresponds to a single predicate-argument relation, expressed in natural language.
@@
Once all QAs are assigned with a label, we can also calculate an overall factual consistency score, defined as the percentage of supported QAs over all QAs. This interpretable score represents the proportion of supported semantic relations within $\mathcal{Y}$ and can be used for model evaluation. For example, the overall factual consistency score of the generated text in Table 1 is 7/12.222In practice, there are a few more QAs that we omitted for brevity. For comparison, the prior localization methods of QGQA and FActScore assign a score of 1/5 to the same summary. This discrepancy stems from the insufficient granularity level of these approaches (see §2.1 and §6), where each assertion contains a mixture of supported and unsupported information.
@@
In this paper, we generate QAs for both verbal and nominal predicates using a parser that we purposely retrained on the QASRL and QANom datasets, improving over the original QASem parser (Klein et al., 2022). Such QAs are generated for each generated text $\mathcal{Y}$, and are then used to detect localized factual inconsistencies by both human annotation (§4) and automated models (§5). See training details and performance of our parser (to be released upon publication) in Appendix A.
@@ 4 Human Detection of Localized Factual Inconsistencies
In this section, we present the QASemConsistency evaluation methodology for manually annotating localized factual inconsistencies in generated texts supported by a corresponding annotation tool (Figure 2), and assess its effectiveness empirically with crowd annotators (§4.1). Our experiments demonstrate the high efficacy of our approach by achieving a high inter-annotator agreement, resulting in an annotation of granular consistency errors (§4.2). We further release a dataset of our annotations to foster future research and evaluation of factual-inconsistency detection models  (§4.3). To the best of our knowledge, this is the first work to annotate factual inconsistencies for predicate-argument level propositions.
4.1 Human Evaluation Process
As mentioned in (§3), given a generated text $\mathcal{Y}$, we predict QA pairs that represent localized propositions corresponding to single predicate-argument relations, using our novel QASem parser. [^3] In some cases, low-quality generated QAs were filtered in the annotation process, in a preliminary step (see Appendix B). Then, human annotators inspect these QAs along with the reference text $\mathcal{X}$ and the generated text $\mathcal{Y}$ and determine for each QA whether it is supported by $\mathcal{X}$ or not. Table 1 (QASemConsistency) shows an example of such annotations. We now describe our complete annotation process.
@@ 4.1 Human Evaluation Process
We first observe that if an entity mentioned in $\mathcal{Y}$ is absent in the reference text $\mathcal{X}$ (i.e., an extrinsic hallucination), all QA pairs featuring this entity as the answer are inevitably not supported. For instance, considering the source article in Table 1, if the summary would have mentioned “An inquest into the death of a woman…”, then the QAs “Who died? A woman” and “Who got something? A woman” would not be supported since the reference text does not mention a woman.
Leveraging this observation, we make the annotation process more efficient by dividing it into two sequential steps. In the first step, annotators go through the entity spans from the generated text $\mathcal{Y}$ corresponding to QASem arguments (i.e. the answers) and classify each span to either “covered” or “not covered”, according to whether it is mentioned explicitly or can be implied from $\mathcal{X}$. Any answer classified as “not covered” (representing an extrinsic hallucination) automatically renders all associated QA pairs as “unsupported”. This eliminates the need for annotators to evaluate these QA pairs individually.
In the second step, annotators focus exclusively on the remaining QA pairs – those whose answers have been confirmed to be covered by the reference text $\mathcal{X}$. We define a QA as “supported” if the meaning of that QA can be deduced from $\mathcal{X}$. To enhance annotators’ focus, all QAs of the same predicate are shown together. In addition to the “support” labels, annotators were encouraged to write free-text notes to justify their decisions, encouraging deeper consideration. These two separate stages improve annotation efficiency, and also introduce an additional layer that classifies factual inconsistencies into extrinsic versus intrinsic errors.
@@
To facilitate the human annotation process, we develop an intuitive annotation interface which streamlines the two steps. [^4] Our tool will be made publicly available. Figure 2 shows the interface of the second annotation step (QA evaluation). See Appendix B for implementation details.
Datasets and Generative Models.
We collect human annotations from two domains and generation settings. First, we consider the task of abstractive summarization on the XSUM dataset, that summarizes news article to a single sentence. Specifically, we sample a subset of 74 summaries from the CLIFF dataset (Cao and Wang, 2021), in which the source articles are from XSUM (Narayan et al., 2018) and the summaries were automatically generated by BART (Lewis et al., 2020) and PEGASUS (Zhang et al., 2020). CLIFF manually annotated each generated summary with token level annotation of consistency errors. Second, we annotate faithfulness localization for 34 generated biographies (3-5 sentences) included in FActScore (Min et al., 2023), which are generated by LLMs such as ChatGPT, InstructGPT, and the retrieval-augmented PerplexityAI model.555perplexity.ai FActScore manually labelled each extracted claim as supported or not against the Wikipedia page of the subject of the biography.
For both datasets, we collect annotations from 3 different workers for each generated text.
@@
We recruit annotators through Amazon Mechanical Turk. [^6] https://www.mturk.com/ Each HIT consists of one reference text, $\mathcal{X}$, and two model outputs, $\mathcal{Y}_{1}$ and $\mathcal{Y}_{2}$. For XSUM, each HIT includes a source article with summaries generated by BART and PEGASUS, and compensated with \$1.50 per HIT. For FActScore, we select two model outputs for each instance, with a compensation of \$3.00 per HIT, reflecting the larger number of QAs to annotate.
Annotation Guidelines.
@@
We follow the controlled crowdsourcing protocol (Roit et al., 2020), which consists of training the workers with detailed instructions (using intuitive slides) and providing ongoing personalized feedback throughout the process. Multiple examples were included to illustrate these two definitions. To further assist annotators with QA evaluation, we advise them to rephrase the question-answer pair as an affirmative statement and assess whether the reference text supports this statement. For instance, the QA pair “What did someone open? An investigation” could be rephrased as “Someone opened an investigation”.
Importantly, in line with our definition of factual consistency as an entailment task (§3), annotators were instructed to rely solely on the reference text. The use of external resources (e.g., web search) was restricted for clarifying the definitions of complex terms (e.g., “six-under-par”), but not to verify content not explicitly stated in the reference text.
4.2 Advantages of QASemConsistency
High Inter-Annotator Agreement.

### v2

@@ Abstract
There has been an increasing interest in detecting hallucinations in model-generated texts, both manually and automatically, at varying levels of granularity. However, most existing methods fail to precisely pinpoint the errors. In this work, we introduce QASemConsistency, a new formalism for localizing factual inconsistencies in attributable text generation, at a fine-grained level. Drawing inspiration from Neo-Davidsonian formal semantics, we propose decomposing the generated text into minimal predicate-argument level propositions, expressed as simple question-answer (QA) pairs, and assess whether each individual QA pair is supported by a trusted reference text. As each QA pair corresponds to a single semantic relation between a predicate and an argument, QASemConsistency effectively localizes the unsupported information. We first demonstrate the effectiveness of the QASemConsistency methodology for human annotation, by collecting crowdsourced annotations of granular consistency errors, while achieving a substantial inter-annotator agreement. This benchmark includes more than 3K instances spanning various tasks of attributable text generation. We also show that QASemConsistency yields factual consistency scores that correlate well with human judgments. Finally, we implement several methods for automatically detecting localized factual inconsistencies, with both supervised entailment models and LLMs. [^1] Our codebase, dataset, and models can be found at https://github.com/ariecattan/qasem_consistency
@@ 1 Introduction
This work focuses on attributable text generation, where the generated content can be verified against a trusted supporting source, referred here as the “reference text”. This reference text may be part of the input for generation, as typical in text summarization or open book QA (Gao et al., 2023b; Slobodkin et al., 2024), retrieved post-generation (Bohnet et al., 2022; Gao et al., 2023a; Min et al., 2023; Wei et al., 2024), or identified by the models themselves when instructed to provide citations to external sources (Liu et al., 2023a; Yue et al., 2023).
@@
In this work, we introduce QASemConsistency, a novel protocol for detecting localized factual inconsistencies in attributable text generation, applicable to both human and automatic detection. Inspired by Neo-Davidsonian formal semantics, our method decomposes the generated text into elementary assertions, in the form of atomic question-answer (QA) pairs (QASRL (He et al., 2015) and QANom (Klein et al., 2020)), where each pair corresponds to a single predicate-argument relation. Localizing factual inconsistencies then involves identifying the set of QA pairs that are not supported by the reference text. Figure 1 illustrates our QASemConsistency methodology by representing the factual inconsistency via a simple QA “Where did someone fall? in the Annalong Valley in County Antrim”, which is not supported by the reference text. By assessing each individual predicate-argument level statement, QASemConsistency can pinpoint more precisely to the factual mistakes than prior methods (illustrated in Table 1). Notably, we found that for 27% of the predicates, some QA pairs were supported and some were not, justifying the need for such a fine-grained representation.
Furthermore, since we represent predicate-argument relations with simple natural language expressions (questions and answers), our QASemConsistency methodology is well-suited for manually annotating granular consistency errors. Indeed, we collect a dataset with localized annotations of factual inconsistency at the predicate-argument level via crowdsourcing and achieve a high inter-annotator agreement (§4). We demonstrate that the overall factual consistency scores obtained by QASemConsistency correlate well with human preferences.
Finally, we implement methods for automatically detecting whether each individual QASem QA is supported by the reference text, following the QASemConsistency methodology (§5). We conduct experiments with a variety of models, including supervised NLI models and prompting open-source and commercial LLMs. While these models were supervised on standard entailment datasets, our results show that they can effectively handle our fine-grained QA assertions, providing more detailed error detection. Yet, there is a vast room for improvement in future work.
Figure 1: An annotation example of a localized factual consistency error according to our QASemConsistency methodology. Here, the model successfully inferred a paraglider fall (after being airlifted with a serious leg injury) but incorrectly identifies the location of the fall as the rescue team’s training area (Annalong Valley) instead of the correct location (Slieve Gullion), which is located 30 miles away. This misattribution is highlighted by the question-answer pair: “Where did someone fall? in the Annalong Valley in County Antrim”.
Altogether, we hope that future research will build upon QASemConsistency for localizing factual inconsistencies (either manually or automatically) and leverage our benchmark to develop better entailment models for handling fine-grained hypotheses corresponding to predicate-argument relations.
@@ 2.2 Predicate-argument Relations: From Neo-Davidsonian Semantics to QA-SRL
$\exists e_{1},e_{2}\,(\text{Buying}(e_{1})\land\text{Agent}(e_{1},\text{Mary})\land\text{Theme}(e_{1},\text{Book})\land\text{Time}(e_{1},\text{Yesterday})\land\text{Giving}(e_{2})\land\text{Agent}(e_{2},\text{Mary})\land\text{Recipient}(e_{2},\text{John})\land\text{Theme}(e_{2},\text{Book})\land\text{Manner}(e_{2},\text{Smile}))$ (1)
@@
To bridge this gap, we propose using QA-SRL, a semantic formalism that simplifies traditional SRL schemes by representing each predicate-argument relation through a simple and “minimal” question-answer pair, such as “Who bought something? Mary” (He et al., 2015; FitzGerald et al., 2018; Roit et al., 2020). For example, the QA-SRL representation of the above sentence is detailed in Table 2. Here, “Mary” is identified as the agent of the predicate “bought”, where this semantic relation is represented by the question “Who bought something? Mary”. Loosely speaking, each QA typically corresponds to a single Neo-Davidsonian proposition.
@@
By relying on a comprehensible natural-language representation, QA-SRL largely subsumes traditional SRL schemes. Notably, it successfully covers valuable implicit semantic arguments (Roit et al., 2024), which are intuitively captured by human annotators, and subsequently by models trained on such annotated data.
@@ 3 QASemConsistency
Given a generated text $y$ and a reference text $x$ that is expected to support the information in $y$ (e.g. the source for generation or a grounding text for it), we define the task of localizing factual inconsistencies as identifying the set of “localized” assertions that are contained in $y$ but are not supported by $x$. This involves the decomposition of $y$ into such assertions, each being a unit of information that can be individually assessed for its entailment by $x$. For effective localization, we suggest two desired properties of this decomposition. First, the assertions should be as minimal in scope as possible, where ideally each assertion should not be further decomposable into smaller verifiable assertions. Second, each assertion should be human interpretable, providing clear insights to common language speakers, thus allowing efficient crowdsourced annotation for localized factual inconsistencies.
To fulfill these properties, we propose decomposing $y$ into the set of its predicate-argument level propositions, using the QASem framework (§2.2). By construction, each question-answer pair (QA) in QASem corresponds to a single predicate-argument relation, expressed in natural language.
@@
Once all QAs are assigned with a label, we can also calculate an overall factual consistency score, defined as the percentage of supported QAs over all QAs. This interpretable score represents the proportion of supported semantic relations within $y$ and can be used for model evaluation. For example, the overall factual consistency score of the generated text in Table 1 is 7/12.222In practice, there are a few more QAs that we omitted for brevity. For comparison, the prior localization methods of QGQA and FActScore assign a score of 1/5 to the same summary. This discrepancy stems from the insufficient granularity level of these approaches (see §2.1 and §6), where each assertion contains a mixture of supported and unsupported information.
@@
We automatically generate QAs for both verbal and nominal predicates using a parser that we purposely trained on the QASRL (FitzGerald et al., 2018) and QANom (Klein et al., 2020) datasets. Specifically, we use the same architecture of the original QASem parser (Klein et al., 2022) and replace the original T5-small model with T5-XL (3B) to improve performance. This parser takes as input a sentence and a predicate, verbal or nominal, and generates a list of atomic QAs. We train our parser for 5 epochs until convergence with the Adam optimizer and a learning rate of 5e055e-05. We evaluate our parser on the QASRL gold data (Roit et al., 2020) and QANom (Klein et al., 2020), achieving 75.9 F1 (+7.3) on QASRL and 72.4 F1 (+13.2) on QANom. Similarly to (Klein et al., 2022), we use a span match threshold of IOU >= 0.3 to match between predicted and gold arguments. The above training datasets do not include annotation of predicate arguments for copular verbs (e.g., “John is a musician”). To ensure completeness of QASemConsistency, we prompt Gemini-Flash (2.0) to generate QAs to represent predicate-argument relations for copular verbs (e.g., “Who is a musician? John”).
@@ 4 Human Detection of Localized Factual Inconsistencies
In this section, we apply our QASemConsistency methodology to manually annotate localized factual inconsistencies in generated texts. We collect such annotations across three different scenarios of attributable text generation: summarization (Cao and Wang, 2021), generation of people’s biography and verification against their Wikipedia pages (Min et al., 2023), and response generation with a generative search engine citing external sources (Liu et al., 2023a).
This annotation serves two primary purposes. First, we assess that QASemConsistency is a suitable approach for collecting high-quality annotations of localized inconsistencies through cost-effective crowdsourcing. This can enable future work to perform human evaluation of generative models with this methodology. Second, we create a diverse and large entailment benchmark where the hypothesis is a predicate-argument relation in the form of a question-answer pair. We believe this benchmark will be valuable for future research to develop and improve models that can predict entailment for predicate-argument level assertions. To the best of our knowledge, this is the first work to annotate factual inconsistencies for predicate-argument level propositions.
4.1 Data Collection
As mentioned in Section 3, given a generated text $y$, we automatically predict QA pairs that represent localized propositions corresponding to single predicate-argument relations, using our QASem parser. [^3] In some cases, low-quality generated QAs were filtered in the annotation process, in a preliminary step (see Appendix A). Then, human annotators inspect these QAs along with the reference text $x$ and the generated text $y$ and determine for each QA whether it is supported by $x$ or not. Table 1 (QASemConsistency) shows an example of such annotations. We now describe our complete annotation process.
@@ 4.1 Data Collection
An entity mentioned in $y$ might be absent in the reference text $x$ (i.e., an extrinsic hallucination) (Xiao and Carenini, 2023). In such cases, all QAs featuring this entity as the answer are inevitably not supported. For instance, considering the source article in Table 1, if the summary would have mentioned “An inquest into the death of a woman…”, then the QAs “Who died? A woman” and “Who got something? A woman” would not be supported since the reference text does not mention any woman.
Leveraging this observation, we make the annotation process more efficient by dividing it into two sequential steps. In the first step, annotators go through the entity spans from the generated text $y$ corresponding to QASem arguments (i.e. the answers) and classify each span to either “covered” or “not covered”, according to whether it is mentioned explicitly or can be implied from $x$. Any answer classified as “not covered” (representing an extrinsic hallucination) automatically renders all associated QA pairs as “unsupported”. This eliminates the need for annotators to evaluate these QA pairs individually.
In the second step, annotators focus exclusively on the remaining QA pairs – those whose answers have been confirmed to be covered by the reference text $x$. We define a QA as “supported” if the meaning of that QA can be inferred from $x$. Specifically, we adhere to the original definition of textual entailment from (Dagan et al., 2013): “a text $T$ entails a hypothesis $H$ if there exists some background knowledge $K$ such that $T$ and $K$ together entails $H$ while $K$ alone does not”. For instance, the reference text “Max was seriously injured when boiling water accidentally spilled on his hand” entails the response “Hot water over 80 degrees Celsius spilled on Max’s hand”, based also on the assumed common knowledge $K$ that “Water boils at 100 degrees Celsius at sea level” but does not entail “Water reaches its boiling point at 100 degrees Celsius”, because the background knowledge $K$ alone suffices to entail the text without the reference text. Therefore, annotators were instructed to rely solely on the reference text $x$ and their common knowledge background to determine whether a QA can be inferred from the reference text. The use of external resources (e.g., web search) was restricted for clarifying the definitions of complex terms (e.g., “six-under-par”), but not to verify content not stated in the reference text.
To further assist annotators with QA evaluation, we advise them to rephrase the question-answer pair as an affirmative statement and assess whether the reference text supports this statement. For instance, the QA pair “What did someone open? An investigation” could be rephrased as “Someone opened an investigation”.
To enhance annotators’ focus, all QAs of the same predicate are shown together. In addition to the “support” labels, annotators were encouraged to write free text notes to justify their decisions, encouraging deeper considerations. These two separate stages improve annotation efficiency and also introduce an additional layer that classifies factual inconsistencies into extrinsic versus intrinsic errors.
@@
To facilitate the human annotation process, we develop an intuitive annotation interface that streamlines the two steps. [^4] Our tool can be found at https://github.com/ariecattan/loc-unfaith. Figure 2 shows the interface of the second annotation step (QA evaluation). See Appendix A for implementation details.
Tasks and Generative Models.
We collect human annotations from three settings of attributable text generation.
First, we consider the task of abstractive summarization on the XSUM dataset that summarizes news articles to a single sentence. Specifically, we sample a subset of 74 summaries from the CLIFF dataset (Cao and Wang, 2021), in which the source articles are from XSUM (Narayan et al., 2018) and the summaries were automatically generated by BART (Lewis et al., 2020) and PEGASUS (Zhang et al., 2020). CLIFF manually annotated each generated summary with token-level annotation of consistency errors.
Second, we annotate faithfulness localization for 34 people biographies (3-5 sentences) included in FActScore (Min et al., 2023). These biographies were generated in zero-shot by LLMs such as ChatGPT, InstructGPT, and the retrieval-augmented PerplexityAI model555perplexity.ai and were subsequently verified against the Wikipedia page of these entities.
Lastly, we extend our evaluation to open-ended response generation with in-line citations to external sources. We use the “Verifiability” dataset (Liu et al., 2023a) that assesses factual consistency of each generated sentence against its corresponding source(s) for several generative search engines (BingChat, NeevaAI, Perplexity.ai, and YouChat) across a range of diverse queries (AllSouls, davincidebate, WikiHowKeywords, ELI5 (KILT / Live), NaturalQuestions). From this dataset, we selected 41 responses, comprising a total of 189 sentences.
Each instance in our dataset includes annotations from 3 different workers.
@@
We recruit annotators through Amazon Mechanical Turk. [^6] https://www.mturk.com/ We follow the controlled crowdsourcing protocol (Roit et al., 2020), which consists of training the workers with detailed instructions (using intuitive slides) and providing ongoing personalized feedback throughout the process. Multiple examples were included to illustrate the two annotation tasks (Span and QA evaluation). Annotators’ compensation is described in Appendix A.
4.2 Dataset Properties
Task #Responses #Sentences #QAs IAA
CLIFF 74 74 693 0.72
FActScore 36 229 1,109 0.78
Verifiability 41 189 1,296 0.67
Table 3: Statistics of our collected benchmark.
Table 3 presents the statistics of our collected dataset. Overall, we gathered entailment annotations for 3,098 different QAs. The ground truth label for each QA is determined by the majority vote among the annotators. We split the dataset into development and test sets 50/50. The development set can serve for prompt engineering or for optimizing the decision threshold.
Each sentence is represented by an average of 6.3 predicate-argument QAs. This granular decomposition contrasts with FactScore’s approach, which yields only 4.1 free-text claims per sentence. In addition, we found that for 27% of the predicates in our collected benchmark, some QA pairs are entailed and some are not. This confirms that assessing each predicate-argument assertions in the model-generated response is valuable and allows to pinpoint the factual inconsistencies.